OpenStudy (anonymous):

W = HR + 1.5VR + B + T, H = 0, B = $300, T = $43.40; find W I got W=0(r)+1.5VR+343.40 What else do i do to solve this

OpenStudy (anonymous):

you almost finish, 0*r=0

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

H*R=0*R=0

OpenStudy (anonymous):

I am confused can you show me what you mean

OpenStudy (anonymous):

i have a total of 8 problems like this i not sure on how to solve can you show me how to solve one

OpenStudy (anonymous):